Solve the problems below and show the complete solution.

What is the f of the pendulum on a clock with a L = 5.00 cm long?
f = (1 / 2π) √ (g / L)
1
Expert's answer
2020-12-08T10:52:23-0500

We can find the frequency of the pendulum from the formula:


"f=\dfrac{1}{2\pi}\sqrt{\dfrac{g}{L}},""f=\dfrac{1}{2\pi}\sqrt{\dfrac{9.8\ \dfrac{m}{s^2}}{0.05\ m}}=2.23\ Hz."

Answer:

"f=2.23\ Hz."